OpenStudy (loser66):

Please, guide me how to convert 623489base 10 to base 16

OpenStudy (anonymous):

Find the biggest power of 16 smaller than 623489 to determine the number of digits and just repeatedly subtract... it's like Euclidean division algorithm

OpenStudy (loser66):

you mean I have to go backward from biggest power of 16, not start from 16^0?

OpenStudy (loser66):

so 16^4 = 65536 then?

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

then?

OpenStudy (phi):

you need a bigger number

OpenStudy (phi):

623489 is much bigger than 16^4

OpenStudy (loser66):

yes

OpenStudy (loser66):

I need more,please, next step?

OpenStudy (phi):

find the n so that 16^n is the first power of 16 bigger than your number then find 16^(n-1) divide 16^(n-1) into your number. the quotient will be the nth digit in your base 16 number. (use A for 10, B for 11, etc if you have to) now do the same thing for the remainder

OpenStudy (loser66):

I got 136*(16^3) +4096 is the next number after 1*(16^4) but how to express 136(16^3)?

OpenStudy (phi):

16^5 is bigger than your number, so 16^4 is what you divide into your number

OpenStudy (phi):

I was under the impression that 623489 was bigger than 16^5, but it is not...

OpenStudy (anonymous):

Yeah phi it's most certainly not

OpenStudy (loser66):

then? how to express it?

OpenStudy (espex):

\(16^0=1\) \(16^1=16\) \(16^2=256\) \(16^3=4096\) \(16^4=65536\) \(16^5=1048576\)

OpenStudy (phi):

how are you getting 1 for the highest digit. what do you get for 623489/16^4

OpenStudy (loser66):

I don't understand your question

OpenStudy (phi):

how many times does 16^4 go into 623489 ?

OpenStudy (e.mccormick):

1 would be the lowest didget. Sounds like you did remainder method backwards.

OpenStudy (loser66):

9

OpenStudy (phi):

the idea is your number is a*16^4 + b* 16^3 + c*16^2 + d*16^1 + e*16^0 a is 9 now subtract 9*16^4 from your number to get the remainder

OpenStudy (espex):

So 623489-(9*16^4) give you the remainder that you move to your next power.

OpenStudy (loser66):

got it, thank you

OpenStudy (loser66):

9(16^4) + 33665 and repeat the step with 33665, right?

OpenStudy (espex):

Right

OpenStudy (loser66):

8(16^3) +... so, if the number before 16^ ... is bigger than 10, for example if it is 13 *(16^2) use A,B ... right?

OpenStudy (phi):

yes,

OpenStudy (loser66):

got it, thanks everybody

OpenStudy (phi):

but you should not be getting any digits > 10 for this problem

OpenStudy (e.mccormick):

Want to see the remainder method for doing this?

OpenStudy (loser66):

yes

OpenStudy (loser66):

I searched but they give me available answer , not method.

OpenStudy (loser66):

my prof taught me calculate from base^0 and up. so I don't know this method

OpenStudy (e.mccormick):

\(\cfrac{623489}{16}=38968\) with a remainder of 1 \(\cfrac{38968}{16}=2435\) with a remainder of 8 \(\cfrac{2435}{16}=152\) with a remainder of 3 \(\cfrac{152}{16}=9\) with a remainder of 8 \(\cfrac{9}{16}=0\) with a remainder of 9 If any remainder was over 9, you would need to convert to the proper letter. Take the remaioners in reverse order: 98381

OpenStudy (loser66):

ok, I know this method, mod(16) and consider the remainder only.

OpenStudy (loser66):

quite easier than go from 16^...

OpenStudy (e.mccormick):

Mathematically, they are related. The remainder method works in reverse, so some people like the power method better because it works the same way they would write out the number.

OpenStudy (espex):

It is also easier to use a calculator with. :)

OpenStudy (loser66):

@eSpeX no calculator, need method.XD

OpenStudy (e.mccormick):

They are pretty close in calculator utility. You have to keep track of fractional values and make them into a useful number either way.

OpenStudy (loser66):

98381 is the answer. I gooooot it hehehe

OpenStudy (loser66):

@e.mccormick you explain me more about the remainder you talked. please

OpenStudy (loser66):

for the first one (remainder 1) the quotient is not an exponent of 16, right? how to link to the problem?

OpenStudy (e.mccormick):

You are dividing out all possible values of 16, the SECOND digit. The remainder of this process is 1, the first digit. So you get 38968.0625 or 38968 with a remainder of 1. That makes 1 the last digit. Now, from 38968 you divide again. This divides out all possible instances of the third digit and leaves a remainder of the second. So on ans do forth. If I used % to mean get the mod, this becomes: \(\cfrac{623489}{16}=38968\) & \(623489\% 16=1\) \(\cfrac{38968}{16}=2435\) & \(38968\% 16=8\) \(\cfrac{2435}{16}=152\) & \(2435\% 16=3\) \(\cfrac{152}{16}=9\) & \(152\% 16=8\) \(\cfrac{9}{16}=0\) & \(9\% 16=9\)

OpenStudy (loser66):

you are wicked!! I "hate" the way you do it!pal

OpenStudy (loser66):

and if the remainder is 10 use A, 11 use B, right?

OpenStudy (e.mccormick):

All I am doing is keepting track of the whole part and the remainder. The whole part goes to the next digit. The remainder goes to the answer. Yes. A=10, B=11, C=12, D=13, E=14, and F=15.

OpenStudy (e.mccormick):

This system also works for base 8, 4, 2, and so on.

OpenStudy (loser66):

thank you